OpenStudy (anonymous):

find the function value of f(x)=[1/((x+4)(x-3))] for f(2)

OpenStudy (anonymous):

1/6

OpenStudy (anonymous):

Because 1/(2+4)(2-3))=1/((6)(1) =1/6